OpenStudy (anonymous):

Use point-slope form of the equation to identify a point the line passes through and the slope of the line. y + 3 = 3(x - 1)

OpenStudy (anonymous):

First solve for Y. y+3=3(x-1), y+3=3x-3, y=3x-6. Your slope is the number in front of the x which is 3 in this case. In order to find a point on the line the easiest way would be to put 0 in for x and solve y= 3(0)-6, y=-6. Therefore, a point is 0, -6